Use the information below to answer the following 2 questions.

A rubber-band slingshot is drawn back is drawn back 0.15 m with a force of 28 N and shoots a 28-g stone. Solve using energy considerations. Recall from the previous lesson that Ep = ½ kx2. Since F = kxEp = ½ Fx.


1)The initial speed of the stone if the rubber band is released is___m/s

2) The maximum height of the stone if it is shot straight up is ___m


Use the information below to answer the following 2 questions.

A 0.80-kg coconut is growing 10 m above the ground in its palm tree. The tree is just at the edge of a cliff that is 15 m tall.


1) The maximum speed of the coconut if it fell to the ground beneath the tree is ___m/s.

2) The maximum speed of the coconut if it fell from the tree to the bottom of the cliff is ___m/s




1
Expert's answer
2021-07-15T12:34:21-0400

"1."

"x = 0.15m"

"F = 28N"

"m= 28g= 0.028kg"

"g= 9.8\frac{m}{s^2}"

"E= E_{r}+E{s}"

"E_{r},E{s}-\text{energy of rubber-band slingshot and stone}"

"\text{before the shot:}"

"E=E_{pr}+E_{kr}+E_{ps}+E_{ks}"

"\text{energy of rubber-band slingshot}"

"E_{kr }=0"

"E_{pr }=\frac{1}{2}Fx=\frac{1}{2}*28*0.15= 2.1 J"

"\text{energy of stone}"

"E_{ks}=0;E_{ps}=0"

"E= 2.1J"

"\text{shot fired: }"

"E=E_{pr}+E_{kr}+E_{ps}+E_{ks} = 2.1J"

"E_{pr}=0;E_{kr}=0;"

"E_{ps}=0;"

"E_{ks}= \frac{mv^2}{2}"

"\frac{mv^2}{2}=E"

"v= \sqrt{\frac{2E}{m}}=\sqrt{\frac{2*2.1}{0.028}}=12.25\frac{m}{s}"

"\text{the stone has reached its highest point:}"

"E=E_{pr}+E_{kr}+E_{ps}+E_{ks}"

"E_{pr}=0;E_{kr}=0"

"E_{ps}= mgh;E_{ks}=0;"

"E = mgh"

"h = \frac{E}{mg}= \frac{2.1}{0.028*9.8}=7.65m"


"\text{Answer:}"

"\text{The initial speed of the stone if the rubber band is released is 12.25 m/s}"

"\text{The maximum height of the stone if it is shot straight up is 7.65 m}"


2.

"\text{Coconut falls from the tree:}"

"h = 10m"

"m = 0.8 kg"

"E= E_p+E_k"

"E_{pmax}= mgh=0.8*9.8*10= 78.4J"

"E_{pmax}= E_{kmax}"

"E_{kmax}=\frac{mv^2}{2}"

"v=\sqrt{\frac{2*E_{kmax}}{m}}=\sqrt{{\frac{2*78.4}{0.8}}}=14\frac{m}{s}"

"\text {Coconut falls from tree and cliff:}"

"h_1 = h+15= 25m"

"E'= E'_p+E'_k"

"E'_{pmax}= mgh_1=0.8*9.8*25= 196J"

"E'_{pmax}= E'_{kmax}"

"E'_{kmax}=\frac{mv_1^2}{2}"

"v_1=\sqrt{\frac{2*E'_{kmax}}{m}}=\sqrt{{\frac{2*196}{0.8}}}=22.16\frac{m}{s}"


"\text{Answer:}"

"\text{The maximum speed of the coconut if it fell to the ground }"

"\text{beneath the tree is 14m/s.}"

"\text{The maximum speed of the coconut if it fell from the tree}"

"\text{to the bottom of the cliff is 22.16m/s}"